Bob employs a painting company that wants to use your services to help create an

invoice generating system. The painters submit an invoice weekly. They charge $30 per

hour worked. They also charge tax – 15%. Bob likes to see an invoice that displays all the

information – number of hours, the charge per hour, the total BEFORE tax and the total

AFTER tax. Since this is an invoice, we also need to display the name of the company

(Painting Xperts) at the top of the invoice. Display all of the information on the screen.

Using Pseudocode, develop an algorithm for this problem.


1
Expert's answer
2021-05-24T07:42:36-0400
start
display (Painting Xperts,Monthly invoce of employees)
working hour per day = x
number of days in a month = z
number of hour in a month y
salary for an hour work = 30$
if(z==30)
  y = x*30
else if(z= 31)
  y = 31*x
else if(z=28)
  y = 28*x
else 
y = 29*x
total_income = y*30
invoice after the tax = total_income - 15*total_income/100
display final invoice(invoice after the tax)